John spends $ 8.40 in buying 100 postage stamps. If x of them are 7 cents stamps & y are 9 cents stamps, find no. of each type of stamps?
Answer by rfer(16322)

x+y=100
x=100-y
------------
0.07x+0.09y=8.40
0.07(100-y)+0.09y=8.40
7-0.07y+0.09y=8.40
0.02y=8.40-7
y=1.40/0.02
y=70
x=100-70
x=30